can you please help me solve this question'
6 2\5 times 2 1\4

Answer by checkley71(8403)

6 2/5=(5*6+2)/5=32/5
2 1/4=(4*2+1)/4=9/4
(32/5)(9/4)=(32*9)/5*4)=288/20=14.4 ANSWER.